# Enhancing Video Large Language Models with Structured Multi-Video Collaborative Reasoning (early version)

### v1

@@ Abstract
Despite the prosperity of the video language model, the current pursuit of comprehensive video reasoning is thwarted by the inherent spatio-temporal incompleteness within individual videos, resulting in hallucinations and inaccuracies. A promising solution is to augment the reasoning performance with multiple related videos. However, video tokens are numerous and contain redundant information, so directly feeding the relevant video data into a large language model to enhance responses could be counterproductive. To address this challenge, we propose a multi-video collaborative framework for video language models. For efficient and flexible video representation, we establish a Video Structuring Module to represent the video’s knowledge as a spatio-temporal graph. Based on the structured video representation, we design the Graph Fusion Module to fuse the structured knowledge and valuable information from related videos into the augmented graph node tokens. Finally, we construct an elaborate multi-video structured prompt to integrate the graph, visual, and textual tokens as the input to the large language model. Extensive experiments substantiate the effectiveness of our framework, showcasing its potential as a promising avenue for advancing video language models.
@@ 1 Introduction
The significant advancement in Large Language Models(LLMs) gives rise to the prosperity of video language models. Equipped with LLMs’ strong general knowledge, video language models showcase an impressive potential to comprehend complex video contents [27, 57, 60, 24, 36, 28, 34]. These models effectively bridge the semantic chasm between video and language, facilitating a nuanced integration of visual perception with language processing.
@@
However, it remains an elusive goal to achieve reliable and comprehensive video reasoning. Specifically, the spatio-temporal knowledge captured from an individual video is often incomplete due to the limited visual perception and comprehending ability. Factors hindering visual perception include sparse sampling [30, 56, 58, 38], spatial occlusion [20, 9], and perspective switching [48, 12, 42, 62]). The comprehending ability is mainly constrained by the complexity of video content [53, 18] and limited effective context length [32, 17]. Since video language models mostly originate from large language models [40, 19] with the help of visual instruction tuning [31], video language models can only resort to the language prior [13, 29] when faced with knowledge incompleteness. This tendency brings spurious hallucination and misalignment with the visual facts [26, 14] to the answers.
@@
The mainstream of video language models follows a single-video reasoning pipeline (shown in Figure 1(a)). Motivated by the research in multi-data collaboration [46, 63, 52] and retrieval-augmented generation [22, 11, 15], a promising idea to address this limitation is integrating knowledge from multiple related videos. We illustrate some video question-answering examples in Figure 2. As the examples demonstrate, single-video reasoning suffers from incomplete and redundant information from an individual video. These challenges prevent video language models from grasping valuable knowledge and lead to hallucinations and failures in answering the given questions. Introducing multiple highly related videos within the same domain can compensate for the missing information and enable the video language models to conclude and provide reliable answers.
@@
Despite the apparent advantages, the incremental information from multiple videos does not necessarily lead to better reasoning in real cases. As illustrated in Figure 1(b), A direct multi-video joint strategy is to concatenate the visual features of multiple videos in the prompt before passing them to the video language models, which can bring an overwhelming number of tokens to video language models [24, 28, 36]. When exposed to long context, empirical evidence [17, 32] proves that LLMs are inclined to concentrate only on specific and limited segments of the input, neglecting potentially critical information. Moreover, the high dimensionality, redundancy, and obfuscation of video content contrast sharply with the structured, clear, and rule-based nature of language. This temporal visual complexity makes the synthesis and refinement of multi-video knowledge particularly challenging.
@@
In this paper, we present a structured multi-video collaborative reasoning framework. As illustrated in Figure 1(c), Our framework achieves the alignment and completion of multi-video information at the graph-structured representation level. Under this framework, to address the redundancy and complexity of video information, we first design the Video Structuring Module (VSM). This module leverages detection and tracking to extract key targets and their spatio-temporal relationships, thereby constructing a data-efficient structured video representation. After obtaining the structured representation, we propose the Graph Fusion Module (GFM) to yield LLM-friendly graph tokens. GFM integrates the structural information into node features through the graph attention network (GAT) and then fuses the multi-video knowledge based on cross-graph attention. Finally, we arrange the fused graph tokens, target video tokens, and text tokens with an elaborately designed prompt to augment VLMs with the multi-video structured knowledge. Overall, our contributions are three-fold:
@@ 2.1 Vision Language Models
In the field of visual understanding, significant advancements have been achieved along with the development of large language models. By learning joint the representation of vision and language through feature alignment and visual instruction tuning [31], vision-language models leverage the generalist capabilities that emerge from the large-scale training of large language models, enabling understanding of images and videos in open-world settings. Among them, Video-LLaMA [60] and BLIP-2 [23] adopt the Q-former to extract valuable visual information into compact and LLM-friendly visual tokens for alignment. On the contrary, Flamingo [1] and BLIP-3 [55] utilize the scalable perceiver resampler to obtain learnable visual tokens. LLaVA [31], Video-LLaVA [28], and Qwen2-VL [43] use an MLP-based projector to map the visual and textual inputs into a shared feature space. Besides, due to the complicated information contained in image and video input, many studies [43, 23, 27] aim to reduce the number of visual tokens for alleviating the computational and understanding burden on LLMs. While current methods focus on single-video reasoning, we try to explore a structured video reasoning framework to effectively address the challenges of redundancy and integration brought by multi-video inputs.
@@ 2.2 Multi-Data Collaboration
Although most deep learning task follows a single data processing pipeline, multi-data collaboration provides a promising direction to improve performances with the inner correspondence among multiple samples. We can roughly divide current works into two streams, which are content-related and task-related collaborations. Content-related collaboration helps models to focus on vital content through comparison among several related data. For example, co-segmentation task [63, 46] explores ways to segment the same object in different scenes by concluding and sharing the same object-related features. Few-shot image classification [16], action recognition [44, 4], and fine-grained classification [35] methods focus on the key differences for accurate classification through multi-data comparison. Retrieval-Augmented Generation [22, 11, 15] is also a promising content-related collaboration that provides the LLMs with supporting information from retrieved related data. In task-related collaboration, models learn the way to complete the task by observing multiple samples performing the same task. For example, multi-video summarization [51, 52, 37] researches aim to generate a summary from a collection of videos through multi-data complementary and refinement. In-Context Learning [45, 3] uses examples with task guidance and answering to demonstrate to LLMs how to fulfill the tasks. For video language models, current multi-video collaboration strategy directly concatenates multiple inputs, leading to burdensome redundancy and difficulties in collaboration. Building on these advancements, we introduce the multi-video reasoning task and achieve the compensation and refinement of key spatio-temporal information through multi-video collaboration.
Figure 3: Multi-video collaborative reasoning framework. Together with the target video, NN related videos are retrieved to facilitate the reasoning process. First, we design the Video Structuring Module to obtain the structured video representation. Then, the Graph Fusion Module fuses the structure information and related videos’ information to get the video graph tokens. Finally, according to designed prompts, the graph tokens, visual tokens and text tokens are arranged as input to the large language model for question answering.
@@ 3.1.1 The setting of multi-video reasoning
In this paper, we explore the potential of utilizing multi-video information for video understanding. Specifically, we follow a multi-video setting, where a target video $V_{0}$ is accompanied by $N$ related videos $\{V_{1},V_{2},\dots,V_{N}\}$. To retrieve related videos, their feature vectors are constructed in advance for efficient video retrieval. Different ways of video vectorization are discussed in supplementary materials. Finally, the methods are required to answer the question to the target video with the help of $N$ retrieved videos.
@@ 3.1.2 Framework overview
Our multi-video collaborative reasoning framework is illustrated in Figure 3. First, the Video Structuring Module is introduced in Sec. 3.2 to get the structured video representation. Based on the obtained video structure, we design the Graph Fusion Module in Sec. 3.3 to fuse the structure information and select useful related videos’ information to the resulting graph tokens. Finally, we arrange all graph, visual, and text tokens according to our designed multi-video reasoning prompt and send them to the large language model.
@@ 3.2 Video Structuring Module
Efficient structured video representation paves the way for the following integration of multi-video knowledge. Given the video and its paired dense caption, the process of our Video Structuring Module (VSM) is depicted on the right bottom of Figure 3, which can be specified as follows.
Step 1: Scene detection. To avoid the video’s temporal redundancy, we use a light-weighted content-based scene detector PySceneDetect [8] to detect scenes from the video. Then, the middle frame of each scene is extracted as the keyframes for the later video structuring process.
@@
Step 2: Dense video captioning. To prepare for the latter structuring pipeline, we need to extract the detailed and fine-grained textual concept. To achieve this goal, we adopt a video language model to give a detailed description of the input video through prompt design, which is detailed in our supplementary materials.
@@
Step 3: Textual scene graph parsing. We then extract the textual scene graph $\mathcal{G}^{\text{Text}}$ from the dense video caption with the help of the SceneGraphParser [50]. The textual scene graph is composed of several triplets $\tau_{i}=\{s_{i},p_{i},o_{i}\}$, where $\tau_{i}$ represents the $i$-th triplets and $s_{i},p_{i},o_{i}$ represents the subject, predicate and object of $\tau_{i}$. Triplets are fundamental objects of interaction and event within the video, formatted in ⟨subject - predicate - object⟩.
Step 4: Graph-guided grounding. Based on the textual scene graph $\mathcal{G}^{\text{Text}}$, we can detect the interacted objects using the open-set detector GroundingDINO [33]. Then, we use the tracking method DEVA [5] to establish the tracklets of the former grounded objects, which temporally connect the same objects. Through detecting and tracking, we can get the: 1) grounding results $\mathcal{B}=\{(b_{s_{i}},b_{o_{i}})|1\leq i\leq M$}, where $b_{s_{i}},b_{o_{i}}$ are the bounding boxes of $s_{i}$ and $o_{i}$, and 2) tracking results $\mathcal{T}=\{\text{Tid}_{s_{i}},\text{Tid}_{o_{i}}\}$, where $\text{Tid}_{s_{i}},\text{Tid}_{o_{i}}$ are the tracking ID of $s_{i}$ and $o_{i}$. Notice that triplets without grounded objects are redundant and should be filtered.
@@
Step 5: Video graph establishment. Built upon $\mathcal{B}$, $\mathcal{T}$ and $\mathcal{G}^{\text{Text}}$, we establish the graph representation for both the target video and related videos. Our graph consists of the nodes, intra-frame edges and inter-frame edges. The node represents the features of the object in the video. To get the node-level features, we use the detecting results $\mathcal{B}$ to crop them from the keyframes and use OpenCLIP [6] to extract their visual features. Specifically, we only utilize the [CLS] token as the output. Since the regions are pre-cropped, the [CLS] token already retains a fine-grained feature representation, eliminating the need to use all output tokens. This significantly reduces the number of tokens required to represent the visual modality. The intra-frame edge is represented by the predicates of each triplet, which indicates the spatial and interacting relationships between objects within one frame. The inter-frame edge connects objects in different frames with the same tracking ID, indicating their temporal relationships. Through the above steps, we can establish the graph-structured representation of both the target video and its related videos for further collaboration.
@@ 3.3 Graph Fusion Module
The Graph Fusion Module (GFM) is a multi-layer stacked architecture (see Figure 3), with each layer comprising two key components: the Hierarchical Frame Graph Attention Network (HF-GAT) [41] and the Cross-Graph Attention (CGA) mechanism.
HF-GAT is tailored to fuse graph-structured data within a single video. Traditional Graph Attention Networks (GATs) are primarily used for tasks such as node classification and relation prediction, where relationships between nodes are explicitly defined. In contrast, for both inter-frame and intra-frame contexts, these relationships are often implicit or absent. To address this challenge, as introduced in Section 3.2, we propose converting the original visual modality data into a graph-structured representation. In this graph, nodes represent features of subjects or objects, which are extracted using OpenCLIP [6]. For intra-frame edges, we utilize relationships based on triplets, where a directional link is formed from $s_{i}$ to $o_{i}$. For inter-frame edges, we leverage the results from the tracker $\mathcal{T}$ to link $s_{i}^{t-1}$ or $o_{i}^{t-1}$ from the previous frame to $s_{i}^{t}$ or $o_{i}^{t}$ in the current frame. Additionally, we introduce bidirectional links for inter-frame connections, as it enhances the system’s ability to comprehend video content [39].
Once the structured features from individual videos are extracted using HF-GAT, the next step is to select the most relevant information from related videos for the target video. To accomplish this, we introduce the cross-graph attention mechanism. Specifically, we use the target video $\mathbf{X}_{tar}$ as the query $\mathbf{Q}_{tar}=\mathbf{X}_{tar}\mathbf{W}_{Q}$, while the concatenation of the target video and related videos along the frame dimension serves as the key and value, i.e. $\mathbf{V}_{tar},\mathbf{V}_{rel}]=[\mathbf{X}_{tar},\mathbf{X}_{rel}]\mathbf{W}_{V}$ and $\mathbf{K}_{tar},\mathbf{K}_{rel}]=[\mathbf{X}_{tar},\mathbf{X}_{rel}]\mathbf{W}_{K}$, where $[,]$ denotes concatenation operation along the frame dimension. The concatenation order of related videos is unordered and interchangeable, so we do not explicitly introduce positional encoding, drawing inspiration from the Set Transformer [21]. The positional encoding within individual videos is implicitly provided by the structured information embedded in the data, as the input to the cross-graph attention mechanism is derived from the output of HF-GAT.
Furthermore, we use class embedding (CE) to enable the cross-graph attention mechanism to distinguish between the graph of the target video and the graphs of related videos.
@@
where $\boldsymbol{\alpha}$ represents a learnable parameter shared across frames. We then apply CE directly to the query, key, and value by adding it to each component:
$\mathbf{\hat{Q}}_{tar}=\mathbf{Q}_{tar}+\mathrm{CE}_{tar},$ (2)
$\mathbf{\hat{K}}_{rel}=\mathbf{K}_{rel}+\mathrm{CE}_{rel},\quad\mathbf{\hat{V}}_{rel}=\mathbf{V}_{rel}+\mathrm{CE}_{rel}.$
Figure 4: Structured multi-video prompts. We properly integrate the multi-modal tokens, together with the one-shot and Chain-of-Thought guidances, to form an LLM-friendly input.
By leveraging CE, the cross-graph attention mechanism implicitly learns to distinguish between the target video and the related videos. The final output of the cross-graph attention is produced by concatenating $\mathbf{O}_{tar}$, the fused representation of the target video, with $\mathbf{X}_{rel}$, the original representation of the related videos that are initially fed into the attention mechanism.
@@
We apply residual connections and pre-norm for both the HF-GAT and cross-graph attention components. Specifically, we use RMSNorm [40], which is commonly employed in large language models (LLMs), for pre-norm. Notably, we omit the Feed-Forward Network (FFN) in this layer to maintain the alignment of features without excessive shifting. The graph-structured representations of the videos are then processed through the GFM to construct graph tokens, each corresponding to the features of subjects or objects after being fused with the structured information.

### v2

@@ Abstract
Despite the prosperity of the video language model, the current pursuit of comprehensive video reasoning is thwarted by the inherent spatio-temporal incompleteness within individual videos, resulting in hallucinations and inaccuracies. A promising solution is to augment the reasoning performance with multiple related videos. However, video tokens are numerous and contain redundant information, so directly feeding the relevant video data into a large language model to enhance responses could be counterproductive. To address this challenge, we propose a multi-video collaborative framework for video language models. For efficient and flexible video representation, we establish a Video Structuring Module to represent the video’s knowledge as a spatio-temporal graph. Based on the structured video representation, we design the Graph Fusion Module to fuse the structured knowledge and valuable information from related videos into the augmented graph node tokens. Finally, we construct an elaborate multi-video structured prompt to integrate the graph, visual, and textual tokens as the input to the large language model. Extensive experiments substantiate the effectiveness of our framework, showcasing its potential as a promising avenue for advancing video language models. Code is open-sourced at https://github.com/ziHoHe/SMV-CR.
keywords:
Structured Multiple Videos, Collaborative Reasoning, Video Structuring Module, Graph Fusion Module, Structured Prompt Engineering
@@ 1 Introduction
The significant advancement in Large Language Models (LLMs) gives rise to the prosperity of video language models (VLMs). Equipped with LLMs’ strong general knowledge, video language models showcase an impressive potential to comprehend complex video contents [li2023llama, yang2023vid2seq, zhang2023video, li2023videochat, maaz2023video, lin2024video, luo2023valley]. These models effectively bridge the semantic chasm between video and language, facilitating a nuanced integration of visual perception with language processing.
@@
However, it remains an elusive goal to achieve reliable and comprehensive video reasoning. Specifically, the spatio-temporal knowledge captured from an individual video is often incomplete due to the limited visual perception and comprehending ability. Factors hindering visual perception include sparse sampling [liu2022task, yang2022tubedetr, yu2021learning, piergiovanni2023rethinking], spatial occlusion [jin2022otpose, dong2023erasing], and perspective switching [shot_switching1, shot_switching2, perspective_shifts1, perspective_shifts2]). The comprehending ability is mainly constrained by the complexity of video content [xia2023structured, ji2020action] and limited effective context length [liu2024lost, hsieh2024ruler]. Since video language models mostly originate from large language models [touvron2023llama, jiang2023mistral, he2025rodimus] with the help of visual instruction tuning [liu2024visual], video language models can only resort to the language prior [goyal2017making, lin2023visualgptscore] when faced with knowledge incompleteness. This tendency brings spurious hallucination and misalignment with the visual facts [li2023evaluating, gunjal2024detecting] to the answers.
@@
The mainstream of video language models follows a single-video reasoning pipeline (shown in Fig. 1(a)). Motivated by the research in multi-data collaboration [wang2023seggpt, zhang2024makes, wu2020dynamic] and retrieval-augmented generation [lewis2020retrieval, edge2024local, guo2024lightrag], a promising idea to address this limitation is integrating knowledge from multiple related videos. We illustrate some video question-answering examples in Fig. 2. As the examples demonstrate, single-video reasoning suffers from incomplete and redundant information from an individual video. These challenges prevent video language models from grasping valuable knowledge and lead to hallucinations and failures in answering the given questions. Introducing multiple highly related videos within the same domain can compensate for the missing information and enable the video language models to conclude and provide reliable answers.
@@
Despite the apparent advantages, the incremental information from multiple videos does not necessarily lead to better reasoning in real cases. As illustrated in Fig. 1(b), A direct multi-video joint strategy is to concatenate the visual features of multiple videos in the prompt before passing them to the video language models, which can bring an overwhelming number of tokens to video language models [li2023videochat, lin2024video, maaz2023video]. When exposed to long context, empirical evidence [hsieh2024ruler, liu2024lost] proves that LLMs are inclined to concentrate only on specific and limited segments of the input, neglecting potentially critical information. Moreover, the high dimensionality, redundancy, and obfuscation of video content contrast sharply with the structured, clear, and rule-based nature of language. This temporal visual complexity makes the synthesis and refinement of multi-video knowledge particularly challenging.
@@
In this paper, we present a structured multi-video collaborative reasoning framework. As illustrated in Fig. 1(c), our framework achieves the alignment and completion of multi-video information at the graph-structured representation level. Under this framework, to address the redundancy and complexity of video information, we first design the Video Structuring Module (VSM). This module analyzes key targets and extracts their spatio-temporal relationships step by step, thereby constructing a data-efficient structured video representation. After obtaining the structured representation, we propose the Graph Fusion Module (GFM) to yield LLM-friendly graph tokens. GFM integrates the structural information into node features through the graph attention network (GAT) [veličković2018graph] and then fuses the multi-video knowledge based on Cross-Graph Attention (CGA). Finally, we arrange the fused graph tokens, target video tokens, and text tokens with an elaborately designed prompt to augment VLMs with the multi-video structured knowledge. Overall, our contributions are threefold:
@@
Figure 3: Multi-video collaborative reasoning framework. Together with the target video, NN related videos are retrieved to facilitate the reasoning process. First, we design the Video Structuring Module to obtain the structured video representation. Then, the Graph Fusion Module fuses the structure information and the related videos’ information to get the video graph tokens. Finally, according to the designed prompts, the graph tokens, visual tokens, and text tokens are arranged as input to the large language model for question answering.
@@ 2.1 Vision Language Models
In the field of visual understanding, significant advancements have been achieved along with the development of large language models. By learning joint representations of vision and language through feature alignment and visual instruction tuning [liu2024visual], vision-language models leverage the generalist capabilities that emerge from the large-scale training of large language models, enabling understanding of images and videos in open-world settings. Among them, Video-LLaMA [zhang2023video] and BLIP-2 [li2023blip] adopt the Q-former to extract valuable visual information into compact and LLM-friendly visual tokens for alignment. On the contrary, Flamingo [alayrac2022flamingo] and BLIP-3 [xue2024xgen] utilize the scalable perceiver resampler to obtain learnable visual tokens. LLaVA [liu2024visual], Video-LLaVA [lin2024video], LLaVA-OneVision [li2024llava], Qwen2/2.5-VL [wang2024qwen2, bai2025qwen2] use an MLP-based projector to map the visual and textual inputs into a shared feature space. Besides, due to the complicated information contained in image and video input, many studies [li2023blip, li2024llama, cheng2024videollama, liu2025nvila] aim to reduce the number of visual tokens for alleviating the computational and understanding burden on LLMs. While current methods focus on single-video reasoning, we try to explore a structured video reasoning framework to effectively address the challenges of redundancy and integration brought by multi-video inputs.
@@ 2.2 Multi-Data Collaboration
Although most deep learning task follows a single data processing pipeline, multi-data collaboration provides a promising direction to improve performance with the inner correspondence among multiple samples. We can roughly divide current works into two streams, which are content-related and task-related collaborations. Content-related collaboration helps models to focus on vital content through comparison among several related data. For example, the co-segmentation task [zhang2024makes, wang2023seggpt] explores ways to segment the same object in different scenes by concluding and sharing the same object-related features. Few-shot image classification [hou2019cross], action recognition [wang2022hybrid, cao2020few], and fine-grained classification [luo2019cross] methods focus on the key differences for accurate classification through multi-data comparison. Retrieval-Augmented Generation [lewis2020retrieval, edge2024local, guo2024lightrag] is also a promising content-related collaboration that provides the LLMs with supporting information from retrieved related data. In task-related collaboration, models learn the way to complete the task by observing multiple samples performing the same task. For example, multi-video summarization [wu2019mvsgcn, wu2020dynamic, panda2017diversity] researches aim to generate a summary from a collection of videos through multi-data complementary and refinement. In-Context Learning [wang2023images, bar2022visual] uses examples with task guidance and answering to demonstrate to LLMs how to fulfill the tasks. For video language models, the current multi-video collaboration strategy directly concatenates multiple inputs, leading to burdensome redundancy and difficulties in collaboration. Building on these advancements, we introduce the multi-video reasoning task and achieve the compensation and refinement of key spatio-temporal information through multi-video collaboration.
@@ 3.1.1 The setting of multi-video reasoning
In this paper, we explore the potential of utilizing multi-video information for video understanding. Specifically, we follow a multi-video setting, where a target video $V_{0}$ is accompanied by $N$ related videos $\{V_{1},V_{2},\dots,V_{N}\}$. To retrieve related videos, their feature vectors are constructed in advance for efficient video retrieval. Different ways of video vectorization are discussed in Sec. 4.4. Finally, the methods are required to answer the question about the target video with the help of $N$ retrieved videos.
@@ 3.1.2 Framework overview
Our multi-video collaborative reasoning framework is illustrated in Fig. 3. First, the Video Structuring Module is introduced in Sec. 3.2 to obtain the structured video representation. Based on the obtained video structure, we design the Graph Fusion Module in Sec. 3.3 to fuse the structured video representation and convert useful related videos’ information to the resulting graph tokens. Finally, we arrange all the graph, visual, and text tokens according to our designed multi-video reasoning prompt and send them to the large language model.
Figure 4: Video captioning prompts. We refer to the design outlined in [zhang2024videoinstructiontuningsynthetic] to create the prompts used to extract captions from videos. The prompts are divided into two parts: the system prompt and the user message. In the system prompt, we define the task of video captioning and provide corresponding guidelines along with a standardized output format. For the output format, the program randomly selects contents in green font as the normalized format for reference during each process of captioning. For the user message, we utilize <<VIDEO_TOKENS>> as the video tokens, and we provide a concise instruction to the model, then generate a detailed description for the video.
Figure 5: Structured multi-video prompts. We properly integrate the multi-modal tokens, together with the prompt guidance, to form an LLM-friendly input.
@@ 3.2 Video Structuring Module
The efficient structured video representation paves the way for the following integration of multi-video knowledge. Given the video and its paired dense caption, the process of our Video Structuring Module (VSM) is depicted on the bottom right of Fig. 3, which can be specified as follows.
Step 1: Scene detection. To reduce temporal redundancy within the video, we employ a lightweight, content-based scene detector, Autoshot [zhu2023autoshot], to segment the video into distinct scenes. From each detected scene, we extract its middle frame as a keyframe, which will serve as input for the subsequent video structuring process. Let the $M$ keyframes of video $V_{N}$ be denoted as $\mathcal{F}_{N}=\{F_{1},F_{2},\dots,F_{M}\}$.
@@
Step 2: Dense video captioning. To prepare for the latter structuring pipeline, we need to extract the detailed and fine-grained textual concept. For this purpose, we leverage a video large language model to generate comprehensive descriptions of the input videos using the designed prompt, as detailed in Fig. 4.
@@
Step 3: Textual scene graph parsing. We then extract the textual scene graph $\mathcal{G}^{\text{Text}}$ from the dense video caption using the SceneGraphParser [wu2019unified], replacing its large language model with Qwen3-30-A3B [yang2025qwen3]. The textual scene graph comprises several triplets $\tau_{i}=\{s_{i},p_{i},o_{i}\}$, where each triplet $\tau_{i}$ represents the $i$-th interaction or event within the video. Here, $s_{i}$, $p_{i}$, and $o_{i}$ denote the subject, predicate, and object, respectively. Each triplet is formatted as ⟨subject - predicate - object⟩, serving as a foundational representation of the relationships and dynamics captured in the video.
Step 4: Graph information filtering. To improve data quality, we apply a proactive filtering mechanism to eliminate irrelevant or redundant triplets. Specifically, we use an image-level classifier to verify the presence of relevant objects or subjects from the triplets within the scene. This is achieved by formulating a simple binary classification task with tailored prompts and utilizing SigLIP [zhai2023sigmoid] to perform the classification. For instance, we design prompts such as, “The object related to {object/subject} is in the image.” for positive samples and, “The object related to {object/subject} is not in the image.” for negative samples. Based on the classification results, we determine whether to retain or discard the object-subject pair. In cases where either the object or subject exists independently in the image, we construct triplets in the format $\{s_{i},*,s_{i}\}$ or $\{o_{i},*,o_{i}\}$, to establish self-connections between nodes in the next step. Conversely, if neither the object nor the subject is present simultaneously, the corresponding triple will be discarded. This process yields the filtered triplets, denoted as $\hat{\mathcal{G}}^{\text{Text}}$.
@@
Step 5: Video graph establishment. Built upon the filtered textual scene graph $\hat{\mathcal{G}}^{\text{Text}}$ and the keyframes $\mathcal{F}_{\{0,\dots,N\}}$ corresponding to each triplet, we establish the graph-based structured video representation for both the target video and related videos. Specifically, the graph consists of the nodes, intra-frame edges, and inter-frame edges. The node represents the features of the object or subject in the video. For node-level features, we first leverage Qwen3-Embedding [zhang2025qwen3] to extract the text features $\mathbf{T}$ from $\hat{\mathcal{G}}^{\text{Text}}$, transforming the text into feature representation for each object and subject, then, we introduce Pooling Attention to extract effective visual features based on text features and the keyframes $\mathcal{F}_{\{0,\dots,N\}}$ and perform adaptive weighted fusion with original text features to obtain a more robust node-level feature representation. Further details regarding this process will be provided in Sec. 3.3. The intra-frame edge is represented by the predicate of each triplet, representing spatial and interacting relationships between objects and objects within one frame. Meanwhile, the inter-frame edge links objects across different frames that share the same subjects and objects, thereby modeling their temporal relationships. Through the above steps, we can establish the graph-based structured video representation of both the target video and its related videos for further collaboration.
@@ 3.3 Graph Fusion Module
The Graph Fusion Module (GFM) consists of a Triplet Embedding Module (TEM) and a multi-layer stacked architecture designed for graph information processing. Each layer of the architecture integrates two essential components: the Hierarchical Frame Graph Attention Network (HF-GAT) and the Cross-Graph Attention (CGA) mechanism.
First, within the TEM, we introduce class embedding (CE) to enhance the ability of the GFM to distinguish between the graph of the target video and the graphs of related videos. The CE is defined as follows:
@@
where $\boldsymbol{\alpha}\in\mathbb{R}^{d}$ represents a learnable parameter shared across frames, and $\sigma$ denotes the sigmoid function. The computed class embeddings are then directly applied to the inputs of the GFM, with $\mathrm{CE}_{tar}$ used for the text features $\mathbf{T}_{tar}$ from the target video and $\mathrm{CE}_{rel}$ for other text features $\mathbf{T}_{rel}$ from the related videos. This integration process is formulated as follows:
$\mathbf{T}_{tar}$ $=\mathbf{T}_{tar}+\mathrm{CE}_{tar},$ (2)
$\mathbf{T}_{rel}$ $=\mathbf{T}_{rel}+\mathrm{CE}_{rel},$
$\mathbf{T}$ $=[\mathbf{T}_{tar},\mathbf{T}_{rel}],$
where $[,]$ indicates the concatenation operation applied to the triplets from the target video and related videos. By leveraging CE, the GFM can implicitly learn to distinguish between the target video and the related videos.
Additionally, to effectively incorporate the visual information from the keyframes corresponding to the triplets, we integrate Pooling Attention, as defined in Eq. (3), within the TEM, enabling direct aggregation of visual features from the keyframes guided by text features.
$\mathbf{Q}$ $=\mathbf{T}\mathbf{W}_{Q}\in\mathbb{R}^{1\times d},$ (3)
$\mathbf{K}$ $=\mathbf{I}\mathbf{W}_{K}\in\mathbb{R}^{(H_{p}\times W_{p})\times d},$
$\mathbf{V}$ $=\mathbf{I}\mathbf{W}_{V}\in\mathbb{R}^{(H_{p}\times W_{p})\times d},$
$\mathbf{\tilde{I}}$ $=\mathrm{softmax}(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d})\mathbf{V}\in\mathbb{R}^{1\times d},$
where $\mathbf{W}_{Q}\in\mathbb{R}^{d\times d},\mathbf{W}_{K}\in\mathbb{R}^{d\times d},\mathbf{W}_{V}\in\mathbb{R}^{d\times d}$ are learnable weight matrices associated with the queries $\mathbf{Q}$, the keys $\mathbf{K}$, and values $\mathbf{V}$ of attention mechanism [vaswani2017attention], respectively, and the $\mathbf{I}\in\mathbb{R}^{(H_{p}\times W_{p})\times d}$ represents the visual features extracted from the vision encoder of VLMs, which utilize multiple tokens to represent a keyframe. $(H_{p}\times W_{p})$ denotes the length of visual features after extraction from the vision encoder. By applying Pooling Attention, we aggregate the visual features in a manner guided by the text features, resulting in a more robust feature representation. Subsequently, we fuse the original text features $\mathbf{T}$, extracted from Qwen3-Embedding [zhang2025qwen3] and the pooled visual features $\mathbf{\tilde{I}}$ using adaptive weights $\boldsymbol{\beta}\in\mathbb{R}^{d}$, as defined by:
$\mathbf{\hat{T}}=\sigma(\boldsymbol{\beta})\odot\mathbf{T}+(1-\sigma(\boldsymbol{\beta}))\odot\mathbf{\tilde{I}}$, (4)
where $\odot$ denotes the Hadamard product (element-wise multiplication). This fusion operation adaptively balances the contributions of text and visual features, yielding a final robust representation $\mathbf{\hat{T}}$.
Then, we input the processed triplet features $\mathbf{\hat{T}}$ into a multi-layer architecture to process the graph information. The features are first passed into HF-GAT, which is tailored to fuse graph-based structured data within a single video. While traditional Graph Attention Networks (GATs) [veličković2018graph] are primarily used for tasks such as node classification and relation prediction, where relationships between nodes are explicitly defined. In contrast, for both inter-frame and intra-frame contexts, these relationships are often implicit or absent. To address this challenge, as introduced in Sec. 3.2, we first convert the original visual modality data into a graph-based structured representation with the VSM. In the constructed graph, nodes represent features of subjects or objects, which are initially extracted from Sec. 3.2 and subsequently processed by the TEM. For intra-frame edges, we utilize relationships based on triplets, where a directional link is formed from $s_{i}$ to $o_{i}$. For inter-frame edges, we leverage the filtered results $\hat{\mathcal{G}}^{\text{Text}}$ obtained in Step 4 (Sec. 3.2) to link $s_{i}^{t-1}$ or $o_{i}^{t-1}$ from the previous frame to $s_{i}^{t}$ or $o_{i}^{t}$ in the current frame. Additionally, we introduce bidirectional links for inter-frame connections, as it enhances the system’s ability to comprehend video content [simonyan2014two].
Once the structured features from individual videos are extracted with HF-GAT, the next step is to identify and fuse the most relevant information between videos. To achieve this, we introduce the Cross-Graph Attention mechanism, implemented through a self-attention mechanism with custom position IDs. To facilitate multi-video collaborative reasoning via triplet features, we identify three key principles: 1) Within a single triplet, the relationship between the subject and object is non-exchangeable. 2) Within a single video, the positional relationship between triplets is unordered. 3) Across multiple videos retrieved via relevance ranking, the order of triplets between videos is non-exchangeable. For principles 1) and 2), the structured video representation from the HF-GAT inherently captures the position encoding within the individual video, as the HF-GAT aggregates and passes information between representations based on the connection relationship of the graph of the corresponding video, thereby implicitly encoding position information in the representation. Consequently, our primary focus is on addressing principle 3) while ensuring compliance with principles 1) and 2). To handle principle 3), we assign consistent position IDs within each video and dynamically adjust position IDs across retrieved videos based on their relevance ranking. For instance, triplet features from the target video are always assigned a position ID of 0, whereas position IDs for related videos are determined dynamically depending on their rank derived from retrieval relevance. These position IDs are then integrated through RoPE [su2024roformer] to encode positional information effectively in the Cross-Graph Attention mechanism.
@@
Moreover, we also apply residual connections and pre-normalization for both the HF-GAT and CGA components. Specifically, we apply the LayerNorm [ba2016layer], commonly used in Vision Transformers (ViT) [zhai2023sigmoid], for pre-normalization. Notably, we exclude the Feed-Forward Network (FFN) in layers to preserve the invariance of aligned visual features from the vision encoder, minimizing excessive feature shifting and ensuring linearity between inputs and outputs of GFM [gao2024linvt]. The graph-based structured video representations of the videos are then processed through the GFM to construct graph tokens, each corresponding to the node features of subjects or objects after being fused with the structured information.